A plane is experiencing a constant force and accel
Answer #1

F = m*a

If we separate this into two differents plane we have the following:

Plane 1: F1 = m1*a1

Plane 2: F2 = m2*a2

However, the Force is constant so F1 = F2, and m2 = 1/2m1, for plane 2 then:

F1 = 1/2m1*a2

2F1/m1 = a2

Then, from the first equation: F1 = m1*a1 ----> m1 = F1/a1, where a1 = 6.5 m/s2 then:

2F1xa1/F1 = a2

a2 = 2a1 = 2*6.5

a2 = 13 m/s2 This would be the correct answer.